Agasse Industries began construction of a new facility and took out a $1,500,000, 8% construction loan on April 1, 2016. Agasse made payments to the general contractor of $481,000 on April 1, $981,000 on August 31, and $581,000 on December 31.

  

Required:
Compute the amount of interest that Agasse would capitalize in 2016. (Do not round intermediate calculations.)

Amount of Interest________
Answer #1
Expenditure
1-Apr 481000*9/9= 481000
31-Aug 981000*4/9= 436000
31-Dec 581000*0/9= 0
917000
interest 917000*8%*9/12= 55020
interest capitalized $55,020